# MIXRTs: Toward Interpretable Multi-Agent Reinforcement Learning via Mixing Recurrent Soft Decision Trees

### v3

@@ 5.1 Environmental description
Spread environment in MPE. We select the commonly used Spread environment, where the objective for a set of $n$ agents is to navigate toward randomly allocated positions marked by $n$ landmarks, concurrently evading any inter-agent collisions. The ideal strategy would involve each agent exclusively claiming a single landmark, thereby achieving an optimal spatial distribution. This necessitates advanced coordination among agents, as they must each predict the targets that their counterparts will likely claim and accordingly adjust their actions to occupy the remaining landmarks. The observation space for a given agent $i$ is represented by a vector that includes the agent’s velocity, its absolute positional coordinates, as well as the relative positions of all other agents and landmarks. These agents possess the capability to traverse in one of four cardinal directions or to remain stationary. The collective reward for the agents is quantified as the negative sum of the minimal distances between each landmark and the nearest agent. To discourage inter-agent collisions, an additional penalizing factor is incorporated into the reward computation, which is defined following the MPE [53]. We consider $3$, $4$, and $5$ agents respectively to complete this experiment and set the training steps to $4M$ to reach convergence.
@@
StarCraft II. This environment is based on StarCraft II unit micro-management tasks (SC2.4.10 version). We consider combat scenarios where the enemy units are controlled by a built-in AI with the difficulty=7 setting, and each allied unit is controlled by the decentralized agents with RL. During battles, the agents seek to maximize the damage dealt to enemy units while minimizing damage received, which requires a range of skills. We evaluate our method on a diversity of challenging combat scenarios, and Table I presents a brief introduction of these scenarios with symmetric/asymmetric agent types and varying agent numbers. To better interpret MIXRTs in such complicated tasks, here we describe the SMAC [33] in detail, including the observations, states, actions, and rewards settings.
@@
Observations and states. At each time step, each agent receives local observations within their field of view, including the following attributes for both allied and enemy units: distance, relative coordination in the horizontal axis (relative X), relative coordination in the vertical axis (relative Y), health, shield, and unit type. All Protos units have shields, which act as a source of defense against attack and have the ability to regenerate if no further damage is dealt. Particularly, the Medivacs as healer units communicate the health status of each agent during the battle. The unit type is maintained for distinguishing different kinds of units on heterogeneous scenarios (e.g., 2s3z and MMM2 in the experiment). Note that the agents can only observe the others if they are alive and within their line of sight range, which is set to $9$. If a unit (for both allies and enemies) feature vector is reset to zeros, it indicates its death or invisibility from another agent’s sight range. The global state is only available to agents during centralized training, which contains information of all units on the map. Finally, all features, including the global state and the observation of the agent, are normalized by their maximum values.
@@
Action space. Each unit takes an action from the discrete action set: no operation (no-op), stop, move [direction], and attack [enemy id]. Agents are allowed to move with a fixed step size in four directions: north, south, east, and west, and the unit is allowed to take the attack [enemy id] action only when the enemy is within its shooting range. Note that a dead unit can only take the no-op action, while a living unit cannot. Lastly, depending on different scenarios, the maximum number of operations that a living agent can perform is typically between $7$ and $70$.
@@
Rewards. The target is to maximize the win rate for each battle scenario. At each time step, the agents receive a shaped reward based on the hit-point damage dealt and enemy units killed. Additionally, agents obtain a $10$ positive bonus after killing each enemy, and a $200$ bonus when killing all enemies, which is consistent with the default reward function of the SMAC. To ensure consistency across different scenarios, the reward values are scaled so that the maximum cumulative reward is around $20$.
@@ 5.2 Experimental Setup
We compare our method to widely investigated value decomposition baselines, including VDN [2], QMIX [3], QTRAN [37], and QPLEX [34], based on an open-source implementation of these algorithms  [^1] https://github.com/oxwhirl/pymarl. Besides, we compare with existing interpretable models of SDTs [27] and CDTs [31] that decompose the problem into a set of simultaneous single-agent problems via the independent Q-learning [44, 54] structure, called as I-SDTs and I-CDTs, respectively. Similarly, for fair comparisons, we also implement RTCs with the independent Q-learning to verify the reliability of the module, called as I-RTCs. The hyperparameters and environmental settings of these algorithms are consistent with their source codes, referring to the SMAC configuration [33]. More details can be found in Table II.
@@
Figure 5: Comparison of I-SDTs, I-CDTs, I-RTCs and MIXRTs with different depths.
During training time, all agents are credited with defeating all enemy units within a limited time after an episode, and the target network is updated once after training every $200$ episodes. We pause the training process for every $5000$ training timesteps and test the win rate of each algorithm for $32$ episodes with the greedy action selection in a decentralized execution manner. All the results are averaged over $8$ runs with different seeds and are displayed in the style of $mean\pm std$. Our model runs from $1$ hour to $15$ hours per task with an NVIDIA RTX 3080TI GPU and an Intel i9-12900k CPU, depending on the complexity and the length of the episode on each scenario.
@@
Figure 6: The effect of MIXRTs on performance with respect to the depth and the number of ensemble trees.
Figure 7: Heatmap visualization of the learned filters and action distributions for each layer. (a) Heatmap visualization of the learned filters in the learned RTCs of depth $3$. The weights of each non-leaf node features contain the current observations (left) and the historical records (right), respectively. The leaf nodes indicate the magnitude of the different action distributions. (b) Actions probability distribution of the nodes of the RTCs with a given observation, where each node is distinguished by a different color bar.
@@ 5.3 Performance Comparison
Performance comparison in MPE. We first run the experiments on three specifically tailored Spread tasks. As shown in Fig. 3, our method achieves competitive performance compared to baselines, demonstrating its efficiency across a range of scenarios. In the standard task configuration involving $3$ agents, most algorithms are successful in learning effective policies that result in covering on average $95\%$ of the landmarks, with the exception of QPLEX. Notably, in scenarios with an increased complexity, featuring $4$ or $5$ agents, MIXRTs demonstrates superior performance in later iterations, achieving a consistent strategy that successfully covers $80\%$ to $90\%$ of the landmarks within $4M$ steps. In contrast, other algorithms display significant performance volatility. This suggests that the deployment of a linear structural tree in policy learning is more conducive to consistently optimal strategy, i.e., each agent is tasked with occupying a distinct landmark.
@@
Performance comparison in SMAC. First, we validate MIXRTs on a range of easy scenarios. As shown in Fig. 4, compared to VDN, QMIX, QTRAN, and QPLEX methods, MIXRTs achieves competitive performance with a slightly faster learning process. On the homogeneous scenarios (e.g., 3m and 8m), MIXRTs performs slightly higher performance during the early learning stage. The baselines obtain a sub-optimal strategy with a slightly larger variance in win rates than the MIXRTs. QTRAN performs not well in these comparative experiments, which may suffer from the relaxations in practice impeding its precise updating [34]. Especially in the heterogeneous 2s3z map, the MIXRTs still obtain a competitive performance whose win percentage is near to QMIX and QPLEX, but slightly higher than VDN and QTRAN, which may benefit from the efficient value decomposition by the lightweight inference.
@@
Next, we test the performance of different algorithms on the hard and super-hard scenarios. Fig. 4 illustrates that MIXRTs consistently achieves performance close to the best baseline on different challenging scenarios, and even exceed it on 6h_vs_8z. For three hard scenarios, MIXRTs obtains slightly superior performance compared to other baselines, underscoring its ability to strike an optimal balance between interpretability and cooperative learning performance. Moreover, we compare the performance of different mixing architectures in these scenarios, as shown in Table IV. The results show that the factorization employed within the mixing trees yields a more effective computation of joint action values than VDN and QTRAN. For super hard tasks 6h_vs_8z and MMM2, MIXRTs can search for workable strategies with stable updates compared to most baselines, indicating that our lightweight inference and ensemble structure can improve the learning efficiency and stability on the asymmetric scenarios. To summarize, MIXRTs achieves competitive performance and stable learning behavior while retaining the concise learning architecture with better interpretability.
@@
In addition, we also analyze the simplicity of MIXRTs in terms of the number of parameters for the above baselines, which depends mainly on the observation size and the number of agents for different scenarios. As shown in Table III, compared to QMIX, QTRAN, and QPLEX, the parameters that need to be learned for VDN are fewer, since there is no mixing network to represent the state-value function. Since each layer of MIXRTs is represented linearly, the number of parameters increases linearly with the depth of the tree. When the depth of the RTCs is set to $3$, compared to QMIX, QTRAN, and QPLEX, the number of parameters of MIXRTs is reduced by more than $49\%$, whereas it retains competitive performance. With the simplicity of the model, we can more easily understand the critical features and the decision process via lightweight inference, which strikes a better balance between performance and interpretability.
Performance comparison with representative interpretable models based on the tree structure. We evaluate the learning performance of RTCs over the individual SDTs and CDTs methods on both easy and hard scenarios. As shown in Fig. 5, I-RTCs constantly and significantly outperforms I-SDTs and I-CDTs in performance and stability for all the tasks. I-RTCs obtains better performance by combining high-dimensional features mapped from observations with historical action-observation values as input, where adding recurrency information could better capture feature information in complex tasks, especially in non-stationary multi-agent tasks. In terms of stability, I-RTCs performs more stably than SDTs and CDTs, its ensemble mechanism can stabilize the training. Compared to the existing interpretable tree-based methods, MIXRTs outperforms I-SDTs, I-CDTs, and I-RTCs as it shows the advantages of learning a centralized but factored centralized joint action-observation value, especially its lightweight inference architecture.
Generally, deeper trees tend to have more parameters and therefore lay more stress on the interpretability. Here, we further analyze the stability of the decision tree methods with different depths on different scenarios. As shown in Fig. 5, I-SDTs, I-CDTs, and I-RTCs learn faster and perform better as the depth of the tree increases. Note that we utilize the ensemble tree structure and the advanced recurrent technique for I-RTCs, yielding better performance and usually being less sensitive to tree depth than I-SDTs and I-CDTs. From the comparisons, MIXRTs yields substantially better results than other trees. The reason might be the effective approximating ability of MIXRTs in the sophisticated relationship between $Q_{i}$ and $Q_{tot}$ with the mixing trees module to capture different features in sub-spaces.
@@
Figure 8: Feature importance and assigned weights of $Q$-values. With the number of steps increases, features importance within the line graph varies correlating to the properties of relevant agents. In the heatmaps, horizontal and vertical ordination indicate the agent-id and steps, respectively.
Sensetivity analysis. In addition, we analyze how sensitive MIXRTs is to the effect of tree depth and the number of ensemble trees on performance. Here, we vary these factors for our experiments to analyze the sensitivity of performance with respect to the tree depth of the individual RTCs and the mixing trees of MIXRTs. Fig. 6(a) and Fig. 6(b) show the respective effects of the above factors on the performance of MIXRTs. First, we investigate the influence of the individual RTCs and the mixing trees with different depths on three scenarios. Fig. 6(a) displays that the performance of MIXRTs will improve as the depth of the individual RTCs and the mixing of RTCs appropriately increases. A greater depth allows MIXRTs to produce more fine-grained behaviors and hence leads to better model performance. However, the improved performance comes at the trade-off of losing its interpretability. Generally, a moderate depth (e.g., depth=3, 3) setting can obtain a competitive performance, where the former $3$ and the latter $3$ represent the depth of the individual action-value RTCs and mixing RTCs of MIXRTs with depth=3$,3$ 3, respectively. Further, we study the effect of the number of ensemble trees on performance. Fig. 6(b) indicates that the tree model shows a clear tendency to be unstable when the number of ensemble trees is small. With moderate values (e.g., $H=32\sim 64$), the tree usually converges quickly and obtains better performance. It shows that the setting with a moderate depth and number of ensemble trees can achieve promising performance while retaining the simplicity and interpretability of the model. For this reason, we chose moderate parameters that are trade-offs between the two, while maintaining a lower amount of parameters comparable to the baselines that can be found in Table III.
@@ 6 Interpretability
The main motivation for this work is to create a model whose behavior is easy to understand, mainly by fully understanding the decision process along the root-to-leaf path and their roles in the team. To display the interpretability afforded by our approach, we describe the structure of RTCs via the learned filters at the inner nodes and present the visualization of the learned probability action distribution. Besides, we present the importance of input features elaborating on how they affect decision making through local features and explore the stability of feature importance. Finally, we give user studies to ensure that the interpretation is consistent with human intuition.
@@
The essence of RTCs is that a kind of model relies on hierarchical decisions instead of hierarchical features. The neural network generally allows the hierarchical features to learn robust and novel representations of the input space, but it will become difficult to interpret once more than one level is present. Compared to the neural network technique, we can immediately engage with each decision made at a higher level of abstraction, where each branching node of the RTCs directly processes the entire input features. By filtering different weights for each feature at each level of branching nodes of the tree, we can understand which features RTCs considers to assign a particular action distribution to a particular state, and how these features influence actions selected by simply examining all the learned filters along the traversed path between the root and the leaf.
@@
As shown in Fig. 7, we display the structure of the learned RTCs model with a depth of $3$ for each agent on a 3m map, where the arrows and lines indicate the connections among tree nodes. Each node assigns different weights to each feature by processing the observed state, where a feature with a more intense color (positive: red, negative: blue) means obtaining a higher magnitude of the weight and receiving more focus. Taking the node in the green dotted area as an example, the three features with more intense color are at position $17$ (representing the relationship with the enemy of $\text{id}=2$), at position $14$ (representing whether this enemy is visible or not), and at position $29$ (representing its own health value). As these positive observations with high weights tend to move to the left leaf, it is essential to select to attack this enemy in the action distribution (marked with a red arrow). Otherwise, the leaf distributions assign probabilities to actions related to moving westward (marked with a green arrow). This behavior is intuitive, as attacking an enemy can bring higher rewards when it has a better chance of survival. Therefore, depending on the choice generated by the activation function, the selected nodes focus on the values of different branching decisions, including action and feature weights.
@@
To understand how a specific state observation influences a particular action distribution, we visualize the decision route from the root to the chosen leaf node in RTCs with the input state. For a given observation, we also provide the action probability distribution for each layer, which is an inherent explanation not provided by the DNNs’ paradigm. As shown in Fig. 7, we can receive the action probability distribution from each layer. Each node outputs a probability distribution with feature vectors, and the selected action depends on the probability linearly combined with the action distribution of each leaf node. From Fig. 7, we can find which features obtain more attention at each layer and how the features affect the action probability distribution. A detailed description is deferred to S. 1 in Supplementary Materials.
@@
TABLE V: Perturbation tests the win rate of MIXRTs. To quantify the interpretation, we use the trained model to calculate the important features in each step via Eq. (15). Then, we mask a varying percentage of least and most important features with zeros for each step and redo the decision making.

### v4

@@ 5.1 Environmental description
Spread environment in MPE. We select the commonly used Spread environment in the MPE, where the objective for a set of $n$ agents is to navigate toward randomly allocated positions marked by $n$ landmarks while evading any inter-agent collisions. The ideal strategy involves each agent exclusively claiming a single landmark, achieving an optimal spatial distribution. This necessitates advanced coordination among agents, as they must predict the targets that their counterparts will likely claim and adjust their actions accordingly to occupy the remaining landmarks. The observation space for a given agent $i$ is represented by a vector that includes the agent’s velocity and absolute positional coordinates, as well as the relative positions of all other agents and landmarks. Each agent can move in one of four cardinal directions or remain stationary. The collective reward for the agents is quantified as the negative sum of the minimal distances between each landmark and the nearest agent. To discourage inter-agent collisions, an additional penalizing factor is incorporated into the reward computation, which is defined following the MPE [53]. We conduct experiments with $3$, $4$, and $5$ agents and set the number of training steps to $4M$ to ensure convergence.
@@
StarCraft II. This environment is based on StarCraft II unit micro-management tasks (SC2.4.10 version). We consider combat scenarios where the enemy units are controlled by a built-in AI with the difficulty=7 setting, and each allied unit is controlled by the decentralized agents using RL. During battles, the agents seek to maximize the damage dealt to enemy units while minimizing damage received, which requires a range of skills. We evaluate our method on a diversity of challenging combat scenarios, and Table I presents a brief introduction to these scenarios with symmetric/asymmetric agent types and varying agent numbers. To better interpret MIXRTs in such complex tasks, we provide a detailed description of SMAC [34] in detail, including the observations, states, actions, and rewards settings.
@@
Observations and states. At each time step, each agent receives local observations within its field of view, including the following attributes for both allied and enemy units: distance, relative coordination in the horizontal axis (relative X), relative coordination in the vertical axis (relative Y), health, shield, and unit type. All Proto’s units possess shields that serve as a defense mechanism against attacks and can regenerate if no further damage is incurred. Notably, Medivacs, being healer units, play a crucial role in maintaining the health of agents during battles. Unit types are used to distinguish different kinds of units on heterogeneous scenarios (e.g., 2s3z and MMM2 in the experiment). Agents can only observe other units if they are alive and within their line of sight range, which is set to $9$. If a unit (for both allies and enemies) feature vector is reset to zeros, it indicates either the unit’s death or its invisibility due to being outside another agent’s sight range. The global state, containing information about all units on the map, is available only to agents during centralized training. Finally, all features, including observations and the global state, are normalized by their maximum values.
@@
Action space. Each unit takes an action from a discrete action set: no operation (no-op), stop, move [direction], and attack [enemy id]. Agents can move with a fixed step size in four directions: north, south, east, and west, and are permitted to execute the attack [enemy id] action only when the enemy is within its shooting range. Note that a dead unit can only take the no-op action, while a living unit cannot select no-op. Finally, depending on different scenarios, the maximum number of operations that a living agent can perform typically ranges from $7$ to $70$.
@@
Rewards. The objective is to maximize the win rate for each battle scenario. At each time step, agents receive a shaped reward based on the hit-point damage dealt and enemy units killed. Additionally, agents receive a bonus of $10$ after killing each enemy, and a $200$ bonus when killing all enemies, which is consistent with the default reward function of the SMAC. To ensure consistency across different scenarios, reward values are scaled so that the maximum cumulative reward is around $20$.
@@ 5.2 Experimental Setup
We compare our method with widely investigated value decomposition baselines, including VDN [2], QMIX [3], QTRAN [38], and QPLEX [35], based on an open-source implementation of these algorithms  [^1] https://github.com/oxwhirl/pymarl. Besides, we compare with existing interpretable models of SDTs [28] and CDTs [32], which decompose the problem into a set of simultaneous single-agent problems via the independent Q-learning [44, 54] structure, referred to as I-SDTs and I-CDTs, respectively. Similarly, to ensure fair comparisons, we also implement RTCs with independent Q-learning to verify the reliability of the module, referred to as I-RTCs. The hyperparameters and environmental settings of these algorithms are consistent with their source codes, adhering to the SMAC configuration [34]. More details can be found in Table II.
@@
Figure 5: Comparison of I-SDTs, I-CDTs, I-RTCs, and MIXRTs with different depths.
[…]
@@
Figure 6: The effect of the depth and the number of ensemble trees on MIXRTs performance.
Figure 7: Heatmap visualization of earned filters and action distributions for each layer. (a) Heatmap visualization of the learned filters in the learned RTCs of depth $3$. The weights of each non-leaf node feature contain the current observations (left) and the historical records (right), respectively. The leaf nodes indicate the magnitude of the different action distributions. (b) Actions probability distribution of the nodes of the RTCs with a given observation, where each node is distinguished by a different color bar.
@@ 5.3 Performance Comparison
Performance comparison in MPE. We first conduct the experiments on three specifically tailored Spread tasks. As shown in Fig. 3, our method achieves competitive performance compared to baselines, demonstrating its efficiency across a range of scenarios. In the standard task configuration involving $3$ agents, all algorithms except QPLEX successfully learn effective policies that result in covering an average of $95\%$ of the landmarks. Notably, in scenarios with increased complexity featuring $4$ or $5$ agents, MIXRTs demonstrate superior performance in later iterations, achieving a consistent strategy that successfully covers $80\%$ to $90\%$ of the landmarks within $4M$ steps. In contrast, other algorithms display significant performance volatility. This suggests that deploying a linear structural tree in policy learning is more conducive to consistently optimal strategy, i.e., each agent is tasked with occupying a distinct landmark.
@@
Performance comparison in SMAC. First, we validate MIXRTs on a range of easy scenarios. As shown in Fig. 4, compared to VDN, QMIX, QTRAN, and QPLEX methods, MIXRTs achieve competitive performance with a slightly faster learning process. In the homogeneous scenarios (e.g., 3m and 8m), MIXRTs perform slightly better than others in the early learning stage. The baselines obtain a sub-optimal strategy with a slightly larger variance in win rates compared to the MIXRTs. QTRAN performs not well in these comparative experiments, which may suffer from the relaxations in practice impeding its precise updating [35]. Especially in the heterogeneous 2s3z map, the MIXRTs still obtain a competitive performance whose win percentage is near to QMIX and QPLEX, but slightly higher than VDN and QTRAN, which may benefit from the efficient value decomposition via the lightweight inference.
@@
Next, we evaluate the performance of different algorithms on the hard and super-hard scenarios. As illustrated in Fig. 4, MIXRTs consistently achieve performance close to the best baseline on different challenging scenarios and even exceed it on 6h_vs_8z. Compared to the other baselines, MIXRTs perform slightly better in the three difficult scenarios, highlighting the model’s ability to strike an optimal balance between interpretability and cooperative learning performance. Moreover, we compare the performance of different mixing architectures in these scenarios, as shown in Table IV. The results show that the factorization employed within the mixing trees yields a more effective computation of joint action values than VDN and QTRAN. For super hard tasks 6h_vs_8z and MMM2, MIXRTs can search for workable strategies with stable updates compared to most baselines, indicating that our lightweight inference and ensemble structure can improve the learning efficiency and stability on the asymmetric scenarios. To summarize, MIXRTs achieve competitive performance and stable learning behavior while retaining an interpretable learning architecture.
@@
In addition, we also analyze the simplicity of MIXRTs in terms of the number of parameters compared to the above baselines, which depends mainly on the observation size and the number of agents on different scenarios. As shown in Table III, VDN requires fewer learnable parameters than QMIX, QTRAN, and QPLEX because it does not employ a mixing network to represent the state-value function. Since each layer of MIXRTs is represented linearly, the number of parameters increases linearly with the depth of the tree. When the depth of the RTCs is set to $3$, compared to QMIX, QTRAN, and QPLEX, the number of parameters of MIXRTs is reduced by more than $49\%$, whereas it retains competitive performance. The simplicity of the model allows us to more easily understand the critical features and the decision-making process through lightweight inference, which strikes a better balance between performance and interpretability.
Performance comparison with representative interpretable models based on the tree structure. We evaluate the learning performance of RTCs over the individual SDTs and CDTs methods on both easy and hard scenarios. As shown in Fig. 5, I-RTCs constantly and significantly outperform I-SDTs and I-CDTs in performance and stability across all the tasks. I-RTCs achieve better performance by combining high-dimensional features mapped from observations to historical information as input, where adding recurrency information can help agents better capture feature information in complex tasks, especially in non-stationary multi-agent tasks. In terms of stability, I-RTCs perform more stably than SDTs and CDTs, owing to its ensemble mechanism. Compared to the existing interpretable tree-based methods, MIXRTs outperform I-SDTs, I-CDTs, and I-RTCs by leveraging the advantages of learning a centralized but factorized joint action-observation value, particularly through their lightweight inference architecture.
Generally, deeper trees tend to have more parameters, which can compromise interpretability. Here, we further analyze the stability of the decision tree methods with different depths across different scenarios. As shown in Fig. 5, I-SDTs, I-CDTs, and I-RTCs learn faster and perform better as the depth of the tree increases. Notably, by utilizing the ensemble tree structure and advanced recurrent techniques in I-RTCs, we achieve better performance that is generally less sensitive to tree depth compared to I-SDTs and I-CDTs. From the comparisons, MIXRTs yield substantially better results than other tree-based methods. This improvement may be attributed to the ability of MIXRTs to efficiently approximate the complex relationship between individual action-values $Q_{i}$ and the joint action-value $Q_{tot}$, as well as to capture different features in subspaces.
@@
Figure 8: Feature importance and assigned weights of $Q$-values. As the number of steps increases, feature importance within the line graph varies correlating to the properties of relevant agents. In the heatmaps, horizontal and vertical ordination indicate the agent-id and steps, respectively.
Sensetivity analysis. In addition, we analyze the sensitivity of MIXRTs to the effects of tree depth and the number of ensemble trees on performance. Here, we vary these factors to assess how performance is influenced by the tree depth of the individual RTCs and the mixing trees of MIXRTs. Fig. 6(a) and Fig. 6(b) show the respective effects of the above factors on the performance of MIXRTs. First, we investigate the influence of the individual RTCs and the mixing trees with different depths in three scenarios. Fig. 6(a) displays that the performance of MIXRTs will improve as the depth of the individual RTCs and the mixing of RTCs appropriately increases. A greater depth allows MIXRTs to produce more fine-grained behaviors, leading to better model performance. However, the improved performance comes at the trade-off of losing its interpretability. Generally, a moderate depth (e.g., depth=3, 3) setting can obtain a competitive performance, where the former $3$ and the latter $3$ represent the depth of the individual action-value RTCs and mixing RTCs of MIXRTs with depth=3$,3$ 3, respectively. Further, we study the effect of the number of ensemble trees on performance. Fig. 6(b) indicates that the tree model tends to be unstable when the number of ensemble trees is small. With moderate values (e.g., $H=32\sim 64$), the tree usually converges quickly and obtains better performance. This suggests that setting a moderate tree depth and number of ensemble trees can yield promising performance while retaining the simplicity and interpretability of the model. For this reason, we chose moderate parameters that offer a trade-off between performance and interpretability, while maintaining a lower number of parameters comparable to the baselines, as shown in Table III.
@@ 6 Interpretability
The main motivation for this work is to create a model whose behavior is easy to understand, mainly by fully understanding the decision process along the root-to-leaf path and their roles in the team. To demonstrate the interpretability provided by our method, we describe the structure of RTCs through the learned filters at inner nodes and show the visualization of the learned action distribution. Furthermore, we present the importance of input features, describe how they influence decision-making, and explore the stability of feature importance. Finally, we give user studies to ensure that the interpretations align with human intuition.
@@
The essence of RTCs is that a kind of model relies on hierarchical decisions rather than hierarchical features. The neural network generally allows the hierarchical features to learn robust and novel representations of the input space, but it will become difficult to interpret once more than one level. In contrast, we can immediately engage with each decision made at a higher level of abstraction, where each branching node of the RTCs directly processes the entire input features. By filtering different weights to each feature at each branching node, it becomes possible to understand which features the RTCs consider when assigning a particular action distribution to a specific state and how these features influence the selected actions. This understanding is achieved by simply examining all the learned filters along the traversed path from the root to the leaf.
@@
As shown in Fig. 7, we display the structure of the learned RTCs model with a depth of $3$ for each agent on a 3m map, where the arrows and lines indicate the connections among tree nodes. Each node assigns different weights to each feature by processing the observed state, where a feature with a more intense color (positive: red, negative: blue) indicates higher magnitude weights and receives more focus. Taking the node in the green dotted area as an example, the three features with more intense color are at position $17$ (representing the relationship with the enemy of $\text{id}=2$), at position $14$ (representing whether this enemy is visible or not), and at position $29$ (representing its health value). Since these features with higher weights tend to direct the decision towards the left leaf, it becomes important to choose to attack this enemy in the action distribution (marked by a red arrow). Otherwise, the leaf distributions assign probabilities to actions related to moving westward (marked by a green arrow). This behavior is intuitive, as attacking an enemy can bring higher rewards when it has a better chance of survival. Therefore, depending on the choice generated by the activation function, the selected nodes focus on the values of different branching decisions, including action and feature weights.
@@
To understand how a specific state observation influences a particular action distribution, we visualize the decision route from the root to the chosen leaf node in RTCs with the input state. For a given observation, we also provide the action probability distribution at each layer, which is an inherent interpretation capability not available by the standard DNNs paradigm. As shown in Fig. 7, we can receive the action probability distribution from each layer. Each node outputs a probability distribution associated with a feature vector, and the selected action depends on the probabilities obtained by linearly combining the action distributions of each leaf node. From Fig. 7, we can find which features obtain more attention at each layer and how these features affect the action probability distribution.
@@
TABLE V: Perturbation tests the win rate of MIXRTs. To quantify the interpretation, we use the trained model to calculate the important features in each step via Eq. (13). Then, we mask a varying percentage of the least and most important features with zeros for each step and redo the decision-making.